OpenStudy (anonymous):

For the system of linear equations below, use graphing to determine in which quadrant the solution lies. x – y = 3 8x + 5y = -2 a.1 b.2 c.3 d.4

OpenStudy (anonymous):

First, solve for y. y=x-3 & y=-8/5x-2/5 Second, graph the lines. (hold on while i draw it) Third, determine where they intersect.

OpenStudy (anonymous):

OMFG I LOVE U!!!! THX FOR ANSWERING!!!

OpenStudy (anonymous):